OpenStudy (anonymous):

Solve. 2 |2x-7| + 11 <- 25

OpenStudy (anonymous):

\[2 |2x-7| + 11 \le 25\]

OpenStudy (nincompoop):

when dealing with absolute values, you end up with two solutions when x is less than zero when x is greater than zero so if for example you have |x-3| it would then be x-3 **** this is greater than zero -(x-3) = -x + 3 **** this is less than zero

OpenStudy (nincompoop):

the rest would be similar to solving any equality or inequality.

OpenStudy (nincompoop):

let me know if you need further clarification

OpenStudy (anonymous):

i dont get it. :O

OpenStudy (anonymous):

so i must multiply the |2x-7| with positive and negative?

OpenStudy (nincompoop):

with positive 1 for one solution and with negative 1 with the other solution

OpenStudy (anonymous):

i see.. let me try to solve it..

OpenStudy (nincompoop):

okay put it here so I can see it

OpenStudy (anonymous):

okay.. here.. \[2 |2x-7| + 11 \le 25\] \[2 +1(2x-7) + 11 \] \[2+2x-7+11 \le 25\] \[2x+6 \le 25\] is this right?

OpenStudy (anonymous):

that's for +1

OpenStudy (nincompoop):

you should avoid using the plus and minus sign, because they can be mistaken for addition and subtraction, which is not the case the 2 outside the | | means to multiply 2 |2x-7| +11 <=25 becomes when x is greater than zero 2 (2x -7) + 11 <=25 when x is less than zero 2 [-1(2x-7)]+11<=25

OpenStudy (anonymous):

ahh.. i see. .okey :D

OpenStudy (nincompoop):

when x is greater than zero, nothing changes, you just copy whatever is inside the absolute bars and proceed with the operation.

OpenStudy (anonymous):

my answer with postive is 4x <_ 27

OpenStudy (anonymous):

and with negative is x <_ 0

OpenStudy (nincompoop):

double check

OpenStudy (anonymous):

okie..

OpenStudy (anonymous):

is it already correct?

OpenStudy (nincompoop):

when x is greater than 0 4x-14+11 <= 25 4x -3 <= 25 4x <= 28

OpenStudy (anonymous):

-14+11= 2

OpenStudy (anonymous):

oopsppss im wrong sorry.

OpenStudy (anonymous):

4x ≤ 28 will i divide both sides by 4?

OpenStudy (nincompoop):

when x is less than zero 2 [-1(2x-7)]+11<=25 2(-2x+7)+11<=25 -4x+14+11<=25 -4x+25<=25 -4x<=0

OpenStudy (anonymous):

can X be a -x?

OpenStudy (nincompoop):

it cannot you are solving for x not -x so when it is -x you multiply it by -1 to undo the negative when it is -4, you divide it by -4 but when dividing an inequality by a negative you will need to flip the inequality from greater than to less than, from less than to greater than, from greater than or equal to to less than or equal to, from less than or equal to to greater than or equal to

OpenStudy (nincompoop):

do you have any lecture notes? these are fundamental that you should have learned a while ago

OpenStudy (anonymous):

ooh.. i see.. then 4x ≤ 28 i should divide both sides by 4?

OpenStudy (nincompoop):

yes

OpenStudy (anonymous):

thx,